OpenStudy (anonymous):

Suppose the function shown below was expressed in standard form, y=ax^2+bx+c. What is the value of a? the graph shows a parabola facing downward with a vertex of (1,4) and the x intercepts are (3,0) and (-1,0) with a y intercept of (0,3)

OpenStudy (anonymous):

vertex tells you it looks like \[y=a(x-1)^2+4\]

OpenStudy (anonymous):

now you find a. by the way you do not need all of this information and it may very well have TOO MUCH INFORMATION.

OpenStudy (anonymous):

the vertex and one other point would do

OpenStudy (anonymous):

for example you know if x = 0 then y = 3 so put \[3=a(0-1)^2+4\] \[3=a+4\] \[a=-1\]

OpenStudy (anonymous):

does that work for any point that you plug in?

OpenStudy (anonymous):

that was enough to find a. you do not need the other points. in fact they could even be wrong. you cannot willy nilly specify 4 points on a parabola

OpenStudy (anonymous):

that is the point! i do not know. we can check them. but you can't just pick 4 points and ask for the parabola. just like i can't say find the equation for a line between 3 points unless they are colinear

OpenStudy (anonymous):

oh alright next time i wont thanks!

OpenStudy (anonymous):

so equation is \[y=-(x-1)^2+4\] and you can check if the other points work or do not

myininaya (myininaya):

there is alot of info here lol

OpenStudy (anonymous):

\[(-1,0)\] works because \[-(-1-1)^2+4=-4+4=0\]

OpenStudy (anonymous):

ikr you can't just say find the parabola that goes through here and here and here and here. maybe there isn't one

OpenStudy (anonymous):

oh i understand so i shouldnt choose random points only ones that i know are correct and then i plug thwem in

OpenStudy (anonymous):

works out in this case because problem was made to work out. but the vertex and one other point would do it

OpenStudy (anonymous):

follow what i did with vertex and one other point above. that should convince you that we had enough information there to find the whole equation. if other points fit that equation then they are on the graph. if they do not then they do not